a housing project has 1822 units. there are twice as many two-bedrooms units as there are one-bedroom units, and 250 fewer three bedroom units than one-bedroom units, how many of each kind of units are there?
Answer by richwmiller(17219)

t+o+r=1822
2o=t
o=t/2
o-250=r
2o+o+o-250=1822
4o-250=1822
4o=1822+250
r=three bed
o=one bed
t=two bed
o=518
t=1036
r=268